OpenStudy (anonymous):

Minimize S = x + 19y with xy = 19 and both x and y > 0. s=? Minimize F = 8x2 + 8y2 with xy2 = 16. F=? Maximize P = xyz with x + y = 3 and y + z = 3, and x, y, and z greater than or equal to 0. P=?

OpenStudy (dumbcow):

these are optimization problems, to minimize or maximize a function, find the point where derivative equals zero for each problem , get function in terms of a single variable, then differentiate and set equal to 0. then solve for value of variable. Use this to get optimized function value. here is how the last one is solved: P = xyz y = 3-x z = 3-y = 3-(3-x) = x therefore, P = x*(3-x)*x = 3x^2 - x^3 differentiate dP/dx = 6x-3x^2 set equal to 0 --> 6x - 3x^2 = 0 --> 3x(2-x) = 0 --> x = 0, x=2 if x=0, then P=0....but that is not a maximum, so use x=2 if x=2, then P = 4 Maximum value of P is 4